OpenStudy (twizttiez):

2. Mendel’s team built a structure for a competition using wooden spheres and cylinders. The wood they used weighs 1 g for each 1.4 cm3. They used 4 spheres and 4 cylinders to build the base of the structure. Each sphere has a radius of 4.45 cm. Each cylinder has a radius of 3 cm. The base weighs 1700 g. What is the height of the cylinders? Use 3.14 to approximate pi and round the answer to the nearest whole number. Show your work.

OpenStudy (anonymous):

ok, lets see here

OpenStudy (anonymous):

haha, this one is a bit tougher

OpenStudy (twizttiez):

Oh no :"""(

OpenStudy (anonymous):

it will take around 2 - 5 minutes

OpenStudy (twizttiez):

Oh that's nothing

OpenStudy (anonymous):

when it says base, is it reffering to whatever these are sitting on top of?

OpenStudy (twizttiez):

I think so

OpenStudy (anonymous):

kk, ill answer based off of that assumption

OpenStudy (twizttiez):

Okaiii

OpenStudy (anonymous):

oohhhh, nvm the question is different than what my thought process was going off of let me rethink this

OpenStudy (twizttiez):

Okay no prob XD

OpenStudy (anonymous):

ok, rounding, it is going to be 51 as the hieght

OpenStudy (anonymous):

here is my work

OpenStudy (anonymous):

ill draw it this time :P

OpenStudy (twizttiez):

Okay

OpenStudy (anonymous):

oh, wait that answer is wrong

OpenStudy (twizttiez):

oooooooo

OpenStudy (anonymous):

its actually 13 -_- silly me, i forgot to divide by 4 before rounding

OpenStudy (anonymous):

now ill show my work

OpenStudy (twizttiez):

Okai

OpenStudy (anonymous):

ok :P i need to rework it

OpenStudy (twizttiez):

Okaii no rush

OpenStudy (anonymous):

ok, i took 4.45*3.14*2*4 = 111.784/1.4 = 79.845714285714285714285714285714 1700 - 79.845714285714285714285714285714 = 1620.1542857142857142857142857143 1620.1542857142857142857142857143 / 3 / 4 = 135.01285714285714285714285714286 that should be rounded to 135 as a whole number for the height of the cylinders, can you tell me were i messed up? i know i messed up somewhere, but its too early in the morning for me to do math properly XD

OpenStudy (anonymous):

i had a red bull late last night, but its wearing off

OpenStudy (twizttiez):

Ehhhhhhhhhhhhhhh I dunno o.o XD

OpenStudy (anonymous):

:P ok one sec, this might take a few more minutes than i thought

OpenStudy (twizttiez):

Okaii

OpenStudy (anonymous):

ok, i found what i did wrong, and im fixing it.

OpenStudy (twizttiez):

Okaii

OpenStudy (anonymous):

ok, finally i got it for sure this time

OpenStudy (twizttiez):

Okay XD

OpenStudy (anonymous):

the technical answer is 5.72, but if you round it is 6

OpenStudy (anonymous):

so i guess 6 would be the whole number they are looking for, and here is how i got it

OpenStudy (anonymous):



OpenStudy (anonymous):

phew, i hope you are satisfied ^_^ that is the answer and how i got it